OpenStudy (anonymous):

Select the correct numeral form of this number written in scientific notation. 5.0 × 10 10 0.000 000 000 5 5,000,000,000 50,000,000,000 500,000,000,000

OpenStudy (e.mccormick):

\(5.0 \times 10^{10}\) OK. So, what is \(10^{10}\)?

OpenStudy (anonymous):

its b 50,000,000,000

OpenStudy (anonymous):

10x10x10x10x10x10x10x10x10x10=10,000,000,000 so 5.0x10,000,000,000=50,000,000,000

geerky42 (geerky42):

\[1\underbrace{0}_{\text{One 0}}=10^1\\~\\1\underbrace{00}_{\text{Two 0s}}=10^2\\~\\1\underbrace{,000}_{\text{Three 0s}}=10^3\\~\\1\underbrace{0,000}_{\text{Four 0s}}=10^4\\~\\1\underbrace{00,000}_{\text{Five 0s}}=10^5\\~\\1\underbrace{000,000 }_{\text{ Six 0s}}=10^6\\~\\~~~~~~~~~~ \vdots\]